Mathematics 243 – Algebraic Structures Definitions Quiz 8 – November 16, 2012

Your Name:

A) (5) Define: G is a group under the operation ∗.

B) (5) Let f : Z → Z_{4} be the mapping that sends each integer x to its congruence class [x] mod 4. Give a description of the set f^{−1}({[0], [3]}).